# Dynamical systems approach to Cold and Warm Inflation within slow-roll and beyond

### v1

@@ 3.2 Case II: $\mathbf{\beta<0}$
It is important to note that from Eq. (3.18), the argument of the inverse sin function must lie within the range $[-1,1]$. This imposes the constraint on $x$:
$-\frac{1}{\sqrt{3}}<x<\frac{1}{\sqrt{3}}\,.$
The above condition shows one of the differences of CRCI, with negative $\beta$, with respect to the case of positive $\beta$. In the present case the range of $x$ is bounded. This also puts a bound on the maximum value of $y$, for $\tilde{\beta}=0.0067$. The bounded interval for $y$ is $(-0.4,0.4)$.
Figure 4: The streamline plots for the field $(\kappa\phi,\dot{\phi}/\sqrt{V_{0}})$ in a specific $(x,y)$ range. The curves correspond to streamlines for $\beta<0$. For these plots $\tilde{\beta}=0.0067$ and the field constant vary. Here green region corresponds to regions where $10.1<\epsilon_{1}<$ 1, pink region corresponds $\epsilon_{1}>1$, and yellow region corresponds to $0.10<\epsilon_{1}<$ 0.1.
@@
4 Warm Inflation (WI)
@@
where $C_{\Upsilon}$ is a dimensionless constant, $c,p$ are the free parameters of the model and $M$ is the mass dimensional constant. In the literature the following forms of the $\Upsilon$ have been studied extensively [37, 39, 38, 40].
$\rm(i)\ T\quad(ii)\ T^{3}\quad(iii)\ \frac{T^{3}}{\phi^{2}}\ .$ (4.6)
The energy density and the pressure of the inflaton field are:
@@
In the subsequent section, we will explore the phase space dynamics corresponding to the quadratic potential of the inflaton field, which takes the following form:
$V(\phi)=\frac{1}{2}m^{2}\phi^{2}\,,$
where $m$ represents the mass of the inflaton field. This potential form remains the same as the form of the potential used in cold inflation.
4.1 Dynamics of the Warm Inflation field
To analyze the dynamics of the inflaton field for various classes of dissipative models described in Eq. (4.6), we define a set of dimensionless dynamical variables to effectively capture the phase space dynamics:
$x=\frac{\kappa\sqrt{V}}{\sqrt{3}H},\quad y=\frac{\kappa\dot{\phi}}{\sqrt{6}H},\quad\Omega_{r}=\frac{\kappa^{2}\rho_{r}}{3H^{2}},\quad z=\frac{H}{m},\quad Q=\frac{\Upsilon}{3H}\ .$ (4.12)
Here, the variables $x$ and $y$ correspond to the field potential and the field’s velocity, respectively, normalized by the Hubble parameter. Additionally, the variable $z$ represents the Hubble parameter $H$, which is introduced to close the system of equations. As we will move over to CRWI in the next section, where we will be only treating cases where $Q$ is a constant or it is a function of temperature, we will like to analyze these two cases in this section so that we can have some comparison of the results. As because we assume thermal stability, we do not separately give detailed analysis about $Q$ depending purely on $T$ as when $T$ is nearly a constant (for thermal stability) $Q$ is also a constant. Consequently, the case of constant $Q$ encompasses both results approximately. Henceforth we present the dynamical analysis assuming a constant $Q$.
The authors in Ref. [53] had initiated a detailed study of the dynamical system related to warm inflation. In the present paper we use the same form of inflaton potential in warm inflation. In this reference the authors worked with a case where $\Upsilon$ was independent of temperature and so was $Q$. In such a case authors did not explicitly use the thermalization constraint and moreover the authors did not construct compactified phase space. As a result of these the authors had to specifically carry out an investigation to figure out the asymptotic solutions. It is assumed that the authors did not test whether $|\dot{T}|/HT<1$ because they were tacitly relying on slow-roll warm inflation where time rate of changes of physical quantities happens slowly. In our case we work with a set of initial conditions and see how the phase space trajectories evolve. We see that there can be trajectories which do not obey the thermal stability criterion (and hence slow-roll criterion) but are dynamically possible. Moreover our phase space dynamics also shows that the slow-roll region does not coincide with the thermally stable region or in other words we can have regions in phase space where $\epsilon_{1}<1$ but $|\dot{T}|/HT>1$.
In Ref. [54] the authors have studied the stability of warm inflationary models which are evolving near the slow-roll limit. They have worked out the conditions of stability of the system near slow-roll evolution. In our work we have focused our attention on the entire available phase space of the system and all possible trajectories. Out of these possible phase space trajectories some correspond to slow-roll evolution. Our work is more focused on the generic phase space morphology of warm inflation and the presentation of the warm inflation dynamics, provided in this section, gives a new way of dealing with warm inflationary stability issues as will become clearer as one goes through the details of the subject presented.
4.1.1 The case where  Q = Constant
The Hubble constraint equation, expressed in terms of the dynamical variables, takes the form:
$1=\Omega_{r}+x^{2}+y^{2}\ ,$ (4.13)
and the Raychaudhuri equation, in terms of the dynamical variables, is expressed as:
$\frac{\dot{H}}{H^{2}}=-3\left(y^{2}+\frac{2}{3}\Omega_{r}\right)\,.$ (4.14)
The other constraints, particularly the thermal stability condition, can be expressed in terms of the dynamical variables as:
$0\leq\Omega_{r}\leq 1,\quad\frac{|\dot{\rho}_{r}|}{H\rho_{r}}=\left|\frac{6Qy^{2}}{\Omega_{r}}-4\right|\ll 1\,.$ (4.15)
The slow-roll parameter, in the present case, becomes
$\epsilon_{1}=3\left(y^{2}+\frac{2}{3}\Omega_{r}\right)\ .$ (4.16)
The autonomous equations corresponding to $(x,y,z)$ can be expressed in terms of the dynamical variables as:
$x^{\prime}$ $=$ $\frac{y}{z}-x\frac{\dot{H}}{H^{2}}\ ,$ (4.17)
$y^{\prime}$ $=$ $-3y(1+Q)-\frac{x}{z}-y\frac{\dot{H}}{H^{2}}\ ,$ (4.18)
$z^{\prime}$ $=$ $z\frac{\dot{H}}{H^{2}}\ .$ (4.19)
Here, $()^{\prime}\equiv d()/dN$, where $N$ is the number of $e$-folds. As the autonomous system of equations requires three independent variables to close the system, the phase space becomes 3-dimensional. From the above equations, it can be observed that the variable $z$ presents an invariant sub-manifold at $z=0$, although the system of equations becomes ill defined near $z=0$. We will tackle the singularity of the system of equations at $z=0$ shortly. According to the definition of $z$ in Eq. (4.12), the value of $z$ is positive. Hence, the region of interest in the phase space is the positive $z$-axis.
Before analyzing the qualitative behavior of the system, let us discuss the range of the dynamical variables to examine the compactness of the phase space. From the Friedmann constraint relation Eq. (4.13), the ranges of the variables $(x,y)$ are:
$-1\leq x\leq 1,\quad-1\leq y\leq 1\ .$ (4.20)
Meanwhile, the range of the variable $z\in[0,\infty)$ makes the phase space non-compact. The value of $z$ near zero implies the value of $H$ is infinitesimally small or $H\ll m$.
To compactify the phase space, we transform the variable $z$ to $\bar{z}$ using the Poincaré transformation:
$\bar{z}=\frac{z}{\sqrt{1+z^{2}}}\ ,$ (4.21)
mapping infinities to a finite range, $\bar{z}\in[0,1)$, transforming the phase space into a cylindrical geometry with unit radius. With this compactification, the new autonomous system of equations becomes:
$x^{\prime}$ $=$ $\frac{y}{\bar{z}}\sqrt{1-\bar{z}^{2}}-x\frac{\dot{H}}{H^{2}}\ ,$ (4.22)
$y^{\prime}$ $=$ $-3y(1+Q)-\frac{x}{\bar{z}}\sqrt{1-\bar{z}^{2}}-y\frac{\dot{H}}{H^{2}}\ ,$ (4.23)
$\bar{z}^{\prime}$ $=$ $(1-\bar{z}^{2})\bar{z}\frac{\dot{H}}{H^{2}}.$ (4.24)
Here, the prime denotes the derivative with respect to the number of $e$-folds, and $\bar{z}$ represents an invariant submanifold at $\bar{z}=0$. However, it must be noted that the autonomous equations diverge as $\bar{z}\to 0$, making this a pathological coordinate. To eliminate this divergence while preserving the qualitative analysis of the system, we adopt a standard procedure commonly used in the dynamical analysis of systems, as discussed in section 2. This method is used in Ref. [15]. The redefined $e$-fold number then becomes:
$dN\mapsto\bar{z}^{n}\ dN^{\prime}.$ (4.25)
The redefinition ensures that the system remains finite at $\bar{z}=0$, validating the system around this critical point. It is important to note that the values of $n$ are model-dependent and must be chosen carefully. These values are determined by analyzing the order of divergence in the dynamical equations. For instance, if the system contains a term of the form $1/\bar{z}^{2}$, the time redefinition should be $dN\mapsto\bar{z}^{2}dN$. Crucially, the redefinition must not exceed the actual order of divergence; overcompensating can inadvertently introduce spurious fixed points and distort the true phase space structure of the system. Additionally, if the expression contains multiple pathological terms—as will be encountered in the subsequent section—these divergences must be addressed collectively, with careful attention paid to the respective orders of divergence to ensure consistent and accurate regularization. Interested readers may refer to Ref. [16] to gain in-depth insight on this technique. In the present case, $n=1$, and Eqs. (4.22)–(4.24) become after redefinition of $N$:
$x^{\prime}$ $=$ $y\sqrt{1-\bar{z}^{2}}-x\bar{z}\frac{\dot{H}}{H^{2}}\ ,$ (4.26)
$y^{\prime}$ $=$ $-3y\bar{z}(1+Q)-x\sqrt{1-\bar{z}^{2}}-y\bar{z}\frac{\dot{H}}{H^{2}}\ ,$ (4.27)
$\bar{z}^{\prime}$ $=$ $(1-\bar{z}^{2})\bar{z}^{2}\frac{\dot{H}}{H^{2}}.$ (4.28)
Fixed points for constant $Q$
Points $(x_{*},y_{*},\bar{z}_{*})$ $\Omega_{\phi}$ $\epsilon_{1}$ $\frac{\dot{\rho}_{r}}{H\rho_{r}}$ Stability
$P_{1}$ $(0,0,0)$ 00 $\sim 2$ $\sim-4$ Stable
$P_{2,3}$ $(\mp 1,0,1)$ $1$ 0 Indeterminate Saddle
$P_{4}$ $(0,0,1)$ 00 $2$ $-4$ Unstable
$P_{5}$ $(0,\sqrt{1+3Q},1)$ $1+3Q$ $3(1+Q)$ $-6(1+Q)$ Saddle
Table 1: The critical points and their nature for WI for constant $Q$.
In all the above equations one has to replace the value of ${\dot{H}}/{H^{2}}$ by the expression of it, given in Eq. (4.14).
Next, we will determine the qualitative behavior of the system by identifying the critical points corresponding to the model. The critical points and their nature for the model are summarized in Tab. [1]. In the table, we have included the stability of the system at these fixed points. The stability is determined by the methods explained initially in section 2 of this article. In the present case we have a constant dissipative factor $Q$. This dynamical system exhibits five critical points, as shown in Tab. [1], where only one fixed point, labeled as $P_{5}$, is $Q$-dependent. At this point, the field fractional energy density is also $Q$-dependent, given by $\Omega_{\phi}=1+3Q$. For any non-zero positive value of $Q$, $\Omega_{\phi}$ exceeds 1, which implies a negative radiation energy density ($\Omega_{r}<0$), rendering this point invalid. Consequently, $P_{5}$ is not considered in the phase space analysis. On the other hand, the remaining fixed points are independent of $Q$.
At point $P_{1}$, all coordinates vanish. This fixed point is unusual, as one will not be able to properly calculate the field energy density or radiation energy density as because the definition of $x,y,\Omega_{r}$ contains division by factors of $H$. As at this point, $z=0$, we have $H=0$. We have shown this fixed point in the plots as because many phase space trajectories travel towards this point. Heuristically, this point represents the far future of the universe when the Hubble parameter tends to zero. This point appears to be a stable fixed point. Near about this point the slow-roll parameter is greater than one showing that this region is not in the inflationary phase and moreover there can be large fluctuations of radiation energy density near about this point.
Points $P_{2}$ and $P_{3}$ represent field-dominated states with a sub-dominant kinetic component ($|y|\ll x$). These points occur in the potential energy dominated phase. In these cases, the slow-roll parameter becomes much smaller than 1, indicating that phase trajectories may come nar to these point during the early inflationary period. However, the thermal condition $\dot{\rho}_{r}/(H\rho_{r})$ cannot be uniquely determined around these points. The system exhibits saddle behavior at these points, characterizing an early accelerating phase.
[…]
Point $P_{4}$ shows unstable behavior and around this point the universe is primarily radiation dominated, scalar field energy density is zero around this point. For the purpose of inflation, this point is not that interesting as is reflectd by the high value of the slow-roll parameter around this point.
The phase space for this model is depicted in Fig. [5] for distinct values of $Q=5$ and $Q=0.5$. The phase space is color-coded, with the light yellow region representing the inflationary phase ($\epsilon_{1}<1$) and the blue region indicating the thermal condition given in Eq. (4.11). The pink lines specify the phase trajectories and from the figures it is clear that the relevant inflationary trajectories are near around the curved wall of the cylindrical phase space.
In the phase space, points $P_{2}$ and $P_{3}$ lie in the yellow region, which overlaps with the blue region. Most trajectories originating from $P_{4}$ initially move toward $P_{2}$ and $P_{3}$ before converging at $P_{1}$, marking the end of the inflationary phase with radiation domination.
Point $P_{5}$ does not appear in the phase space since any positive $Q$ results in a negative radiation energy density, making it physically non-viable. The behavior of trajectories within this confined phase space remains consistent. It is seen that for lower values of the dissipative factor, a greater number of trajectories converge to $P_{1}$.
The evolution of cosmological parameters, including the slow-roll parameter, thermal stability condition, and fractional energy densities, is shown in Fig. [6] as a function of the redefined number of e-folds $N^{\prime}$, based on selected initial conditions. In the very early phase, up to $N^{\prime}\sim 55$, the field energy density dominates over the radiation energy density. During this epoch, the slow-roll parameter remains less than 1, indicating that the field drives an accelerating inflationary phase and $T>H$.
Throughout this phase, the system is nearly thermally stable. This is illustrated by the modulus of $\left|\frac{\dot{\rho}_{r}}{H\rho_{r}}\right|$, which remains below 1 during the inflationary regime, confirming the thermal stability of the system. In Fig. [6] we have also plotted the original $e$-folds with the redefined $N^{\prime}$. This plot shows that the system comes out of inflationary regime before 60 $e$-folds. This information by itself is not a difficult problem as we have worked with some parameters and with some initial condition to generate this particular dynamical flow. Changing the initial conditions one can always produce the desired number of $e$-folds in the present case. The plots furnished here shows that our dynamical systems approach can reproduce all the essential calculations relating to warm inflation.
Figure 6: The evolution of cosmological variables for the case where $Q=5.0$ and the plot of $N$ with respect to $N^{\prime}$. For the entire inflationary regime we have checked that $T/H>1$.
The analysis of the constant $Q$ warm inflation shows that in this case also we can show the nature of the trajectories in the three dimensional phase space of the system. The compactified phase space show cylindrical symmetry and do show that many trajectories do get attracted towards an inflationary regime. In this section we have introduced all the relevant methods in which one can study the stability of warm inflationary systems of many more involved systems. Next we will try to see how the phase space dynamics of warm inflation in the constant $Q$ regime is modified when we apply the constant-roll condition. One must note that the modification will not in general correspond in a simple way to the unmodified case, presented here, because the form of the potential giving rise to CRWI will change from the one used here. Inspite of the different forms of the potentials we will be able to figure out how the new dynamics in CRWI differs from simple warm inflation dynamics.
5 Constant-roll warm inflation (CRWI)
@@ 5.1 CRWI with $\beta>0$ and constant $Q$
One can see that the inflaton potential used in the present context is not positive definite but for inflation we will be working with a class of initial conditions and a particular set of parameter values for which $V(\phi)>0$. As because $V$ can be negative, we will have to work only with the positive or negative branch of $x$. Here we work with positive values of $x$. The primary variables $(x,z)$ can take any values within the range $[0,\infty)$. To keep track of the behavior of the variables at infinity we compactify the space spanned by these variables and define:
@@
$\bar{x}^{\prime}$ $=$ $-{y}\mathcal{A}\sqrt{(1+Q)\beta}\dfrac{(1-\bar{x}^{2})^{2}\sqrt{\frac{1}{\mathcal{A}}\dfrac{2\bar{x}^{2}+1}{1-\bar{x}^{2}}}\sqrt{\frac{1}{\mathcal{A}}\dfrac{2\bar{x}^{2}+1}{1-\bar{x}^{2}}-1}}{\bar{x}}\ ,$ (5.8)
@@
$P_{1}$ $(0,0)$infty $\infty$ $-\frac{3Q}{2}-1$ Unstable
$P_{2}$ $(1,1)$ Indeterminate $2$ $-1$ Unstable
$P_{3}$ $(\rm{Any},1)$ 00 $2$ $-1$ Unstable
@@
$\bar{x}^{\prime}$ $=$ $-{y}\bar{z}^{2}\mathcal{A}\sqrt{(1+Q)\beta}{(1-\bar{x}^{2})^{4}\sqrt{\frac{1}{\mathcal{A}}\dfrac{2\bar{x}^{2}+1}{1-\bar{x}^{2}}}\sqrt{\frac{1}{\mathcal{A}}\dfrac{2\bar{x}^{2}+1}{1-\bar{x}^{2}}-1}}\ ,$ (5.12)
$\bar{z}^{\prime}$ $=$ $\bar{z}^{2}\bar{x}\left(1-\bar{z}^{2}\right)^{\frac{3}{2}}\left(-y^{2}-2\dfrac{\bar{z}^{2}}{1-\bar{z}^{2}}+2\dfrac{\bar{x}^{2}}{1-\bar{x}^{2}}\right)(1-\bar{x}^{2})^{2}\ .$ (5.13)
@@
Figure 8: The phase space is corresponding to CRWI when $\beta>0$. The color scheme highlights the different dynamical regions: pink ($\epsilon_{1}\geq 1$), yellow region ($0\leq\epsilon_{1}\leq 0.1$), green region ($10.1<\epsilon_{1}<$ 1), blue (thermally stable region), and violet (overlap of red and blue regions).
The phase space is depicted in Fig. [8] for different values of $(Q,\beta)$ in the parameter space of $(\bar{x},\bar{z})$. The phase space is illustrated using distinct color regions: pink represents $\epsilon_{1}\geq 1$, green corresponds to $10.1<\epsilon_{1}<$ 1, yellow indicates $0\leq\epsilon_{1}\leq 0.1$, and violet denotes the thermally stable region. The white region represents an unphysical domain where $\epsilon_{1}$ becomes negative.
@@
[…]

### v3

@@ 2.2 Case II: $\mathbf{\beta<0}$
It is important to note that from Eq. (2.24), the argument of the inverse sin function must lie within the range [1,1][-1,1]. This imposes the constraint on $x$:
$\frac{1}{\sqrt{3}}<x<\frac{1}{\sqrt{3}}\,.$
The above condition shows one of the differences of CRCI, with negative $\beta$, with respect to the case of positive $\beta$. In the present case the range of $x$ is bounded. This also puts a bound on the maximum value of $y$, for $\tilde{\beta}=0.0067$. The bounded interval for $y$ is (0.4,0.4)(-0.4,0.4).
Figure 3: The streamline plots for the field $(\kappa\phi,\dot{\phi}/\sqrt{V_{0}})$ in a specific $(x,y)$ range. The curves correspond to streamlines for $\beta<0$. For these plots $\tilde{\beta}=0.0067$ and the field constant vary. Here green region corresponds to regions where $0.1<\epsilon_{1}<1$, pink region corresponds $\epsilon_{1}>1$, and yellow region corresponds to $0<\epsilon_{1}<0.1$.
@@
The results presented in this section can be compared with the result of another contemporary work in this field. In Section 2.5 of Ref. [73], the authors state that for the range ${3}/{2}<\beta<0$, the background solution is an attractor. Since we have worked with $\beta=-0.0067$, which lies within this interval, our results are fully consistent with their analysis: we also obtain an attractor solution. Moreover , in Section 2.4 of the above reference, one of the values of $\beta$ considered by the authors is $0.01$, which is in close proximity to one of the values we have used in the previous subsection, namely $\beta=0.009$. These facts show that our parameter choices are consistent with the latest work on this field.
3 Constant-roll warm inflation (CRWI)
@@
where $C_{\Upsilon}$ is a dimensionless constant, $c,p$ are the free parameters of the model and $M$ is the mass dimensional constant. In the literature the following forms of the $\Upsilon$ have been studied extensively [53, 10, 13, 18]. The energy density and the pressure of the inflaton field are:
@@ 3.1 CRWI with $\beta>0$ and constant $Q$
One can see that the inflaton potential used in the present context is not positive definite but for inflation we will be working with a class of initial conditions and a particular set of parameter values for which $V(\phi)>0$. As because $V$ can be negative, we will have to work only with the positive or negative branch of $x$. Here we work with positive values of $x$. The primary variables $(x,z)$ can take any values within the range $0,\infty)$. To keep track of the behavior of the variables at infinity we compactify the space spanned by these variables and define:
@@
$\bar{x}^{\prime}$ $=$ $-\dfrac{(1-\bar{x}^{2})^{2}\sqrt{\frac{1}{\mathcal{A}}\dfrac{2\bar{x}^{2}+1}{1-\bar{x}^{2}}}\sqrt{\frac{1}{\mathcal{A}}\dfrac{2\bar{x}^{2}+1}{1-\bar{x}^{2}}-1}}{\bar{x}}\times$ (3.18)
$\left({y}\mathcal{A}\sqrt{(1+Q)\beta}\right)\ ,$
@@
$P_{1}$ $(0,0)$ $\infty$ $\infty$ $\frac{3Q}{2}-1$ Unstable
$P_{2}$ $(1,1)$ Indeterminate 22 1-1 Unstable
$P_{3}$ $(\rm{Any},1)$ 0 22 1-1 Unstable
@@
$\bar{x}^{\prime}$ $=$ $-\bigg({y}\bar{z}^{2}\mathcal{A}\sqrt{(1+Q)\beta}(1-\bar{x}^{2})^{4}\sqrt{\frac{1}{\mathcal{A}}\dfrac{2\bar{x}^{2}+1}{1-\bar{x}^{2}}}\sqrt{\frac{1}{\mathcal{A}}}$ (3.22)
$\dfrac{2\bar{x}^{2}+1}{1-\bar{x}^{2}}\bigg)-1\ ,$
$\bar{z}^{\prime}$ $=$ $\bar{z}^{2}\bar{x}\left(1-\bar{z}^{2}\right)^{\frac{3}{2}}\left(-y^{2}-2\dfrac{\bar{z}^{2}}{1-\bar{z}^{2}}+2\dfrac{\bar{x}^{2}}{1-\bar{x}^{2}}\right)$
$\times(1-\bar{x}^{2})^{2}\ .$
@@
Figure 5: The phase space is corresponding to CRWI when $\beta>0$. The color scheme highlights the different dynamical regions: pink ($\epsilon_{1}\geq 1$), yellow region ($0\leq\epsilon_{1}\leq 0.1$), green region ($0.1<\epsilon_{1}<1$), blue (thermally stable region), and violet (overlap of red and blue regions).
The phase space is depicted in Fig. [5] for different values of $(Q,\beta)$ in the parameter space of $(\bar{x},\bar{z})$. The phase space is illustrated using distinct color regions: pink represents $\epsilon_{1}\geq 1$, green corresponds to $0.1<\epsilon_{1}<1$, yellow indicates $0\leq\epsilon_{1}\leq 0.1$, and violet denotes the thermally stable region. The white region represents an unphysical domain where $\epsilon_{1}$ becomes negative.
@@
As the radiation energy density starts to dominate, the slow-roll parameter exceeds 1, signaling the system’s exit from the inflationary phase after ${t^{\prime}}=195$ which corresponds to about 87 $e$-folds. It must be noted that by tuning the initial conditions one can reduce the number of $e$-folds. This tuning requires a detailed search in the initial condition space and parameter space of the theory. In this paper we present the method in principle and show that a through numerical stability analysis is possible in highly constrained inflationary systems. To calculate the number of e-folds, we use the formula: $N=\int H\,dt,$ which, in our case, can be expressed as:
$N=\int H\,dt=\int\kappa\sqrt{V_{0}}z\,dt=\int\frac{\bar{z}}{\sqrt{1-\bar{z}^{2}}}\,dt^{\prime}.$